# A General Class of Model-Free Dense Precision Matrix Estimators

### v1

@@ 4 Uniform Consistency of General Estimator
In this section we provide first upper bounds on $l_{2}$ norm errors for each regression coefficient estimate, that is in (11). Next, we consider (12), and obtain upper bounds for the estimation errors. Analysis of (12) is key and we provide a new proof technique with using hidden factors without any exact sparsity assumptions. We use concentration inequalities for spectral norms, and high dimensional probability concepts to obtain upper bounds. Then we consider the consistency of each row of precision matrix estimate.
@@ 4.1 Bounding Estimation Errors: Regression Coefficients
which is the $\hat{r}_{j}$- th largest singular value (squared, scaled by n) of $Y_{-j}P_{\hat{B}}$ ($n\times p-1$ matrix). This last term is obtained when we project $Y_{-j}$ onto range of $\hat{B}$, and will reduce the bias of estimating the rows of the precision matrix in Theorem 1 below. Next, we define
@@
Define $\max_{1\leq j\leq p}\hat{r}_{j}:=\bar{r}$, $\min_{1\leq j\leq p}\hat{\eta}_{j}:=\bar{\eta}$, and $\max_{1\leq j\leq p}\hat{\Psi}_{j}:=\bar{\Psi}$. We have the following uniform bounds for estimators for off-diagonal terms in each row of the precision matrix
Theorem 1.
@@
$\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\sigma^{2}\left(\frac{2logn+\bar{r}}{n\bar{\eta}}\right)+\frac{CK}{\bar{\xi}}\left(1+\frac{1}{\delta_{n}}+\frac{\bar{\Psi}}{\bar{\eta}\delta_{n}}+\frac{r_{n}}{\bar{\eta}\delta_{n}}\right).$
(ii) Under Assumptions 1-4, and $r_{n}/\delta_{n}\to 1$,
$\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}^{2}=O_{p}\left(\max(\frac{K}{\bar{\xi}},\frac{K\bar{\Psi}}{\bar{\xi}\bar{\eta}\delta_{n}},\frac{K}{\bar{\xi}\bar{\eta}},\frac{logn+\bar{r}}{n\bar{\eta}})\right)=O_{p}(r_{w2}^{2}),$
where $r_{w2}$ is defined in (9)-(10), by using $\delta_{n}$ is nondecreasing in $n$.
Remark 2.
  • One key issue is that our Theorem 1 forms a key building block for the consistency of rows of precision matrix estimation. The proof of this theorem uses a uniform concentration inequality type result that we show through a novel approach.
  • The first term on the right side of Theorem 1(i) is the variance term and the second term is the bias term.
  • There is no sparsity parameter on the upper bounds and rates.
  • Note that $\bar{\xi}$ plays an essential role in getting consistency. Larger signal-to-noise ratio is needed, especially if the number of factors $K$ grow fast.
  • The upper bounds given in the theorem are general. They do simplify as one moves to particular analyses of RRE and adaptive PCR subcases. However, it is useful to point out some generic principles. For example, $\bar{r}$ represents complexity, and affects the $l_{2}$ norm error adversely through the increase of the estimator’s variance. On the other hand, $\bar{\eta}$ is a positive influence, since a better prediction of asset returns gives a lower $l_{2}$ error. It should particularly be noted that this positive influence is reflected through the reduction of both the estimator’s bias and variance. $\bar{\Psi}$ shows how badly hidden factor estimates predict asset returns and it has an adverse effect on $l_{2}$ error. This term increases the bias of the estimator.
@@ 4.1.1 Adaptive PCR
where $(.)^{+}$ is Moore-Penrose inverse of a matrix.
Before simplifying the terms in the upper bound in Theorem 1 we define $\hat{\lambda}_{k_{j}}:=\frac{1}{n}\sigma_{k_{j}}^{2}(Y_{-j})$, for example with $k_{j}=1$ we have the largest singular value of $Y_{-j}$, which is $\sigma_{1}^{2}(Y_{-j})/n$ (squared-scaled by $n$). One big issue is the choice of $k_{j}$. To choose $k_{j}$ we adapt the so-called elbow method, (for details, see (19) of Bing et al. (2021))
@@
We show in the proof of Corollary 1 in Online Appendix
@@
which will simplify the upper bound for $l_{2}$ norm error for our estimators. We prove the following error bounds for adaptive PCR based estimator for the off-diagonal terms in a given row for the precision matrix.
Corollary 1.
Under Assumptions 1-4 with $10<c^{\prime}<$ 1 a positive constant, $Klogn\leq cn$, with $c$ a positive constant, and $r_{n}/\delta_{n}\to 1$
@@
$\max_{1\leq j\leq p}\|\hat{\alpha}_{PCR,j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\sigma^{2}\left[\frac{2logn+K}{n\delta_{n}}\right]+\frac{CK}{\bar{\xi}}\left[1+\frac{3}{\delta_{n}}\right],$
@@
$\max_{1\leq j\leq p}\|\hat{\alpha}_{PCR,j}-\alpha_{j}^{*}\|_{2}^{2}=O_{p}\left(max(\frac{logn+K}{n\delta_{n}},\frac{K}{\bar{\xi}})\right).$
Remark 3.
  • This is a new result for precision matrix row estimation. It clearly shows, that the estimator’s variance (i.e. the first square bracketed term on (i)-(ii)) goes to zero if $(logn+K)/n\delta_{n}\to 0$. Given $Klogn<cn$ is assumption, this is easily satisfied. It should be noted that large number of factors can be detrimental unless noise $\delta_{n}$ is large.
  • Regarding the bias term (the second square bracket in (i)) and the corresponding rate in (ii) the signal to noise ratio $\bar{\xi}$ should grow larger than $K$ to have consistency of the estimator. Utilizing, for example (8), we may need $Kr_{n}/d_{1n}^{2}\to 0$.
  • We note that with $p>n$, consistent estimation of $\alpha_{j}^{*}$ is possible, and this is also true with $p\leq n$ as well.
  • If we fix $p/n=\gamma\in[0,\infty)$ and then we increase $n,p$ together, and we keep $\gamma$ fixed to check consistency. In $(logn+K)/(n\delta_{n})$ condition, $\delta_{n}$ can grow with $n,p$ both increasing and have a positive effect on consistency. The price here is $logn$ rate, but that is not large compared to $n$ rate in the denominator. In the second condition, $K/\bar{\xi}$ is via (8) tied to $Kr_{n}/d_{1n}^{2}$ rate as discussed in second item in Remark 3 here. Increasing $n,p$ together can increase $r_{n}$, but also $d_{1n}$. Clearly, one observes a tradeoff in the $\delta_{n},r_{n}$ (both representing the noise) role in getting consistency. In particular, the noise can reduce variance but can increase bias. Note that $d_{1n}$ shows the strength of factor loadings $A_{-j}$. With $Kr_{n}/d_{1n}^{2}\to 0$ we have consistency.
[…]
@@ 4.1.2 Ridgeless Regression Estimation
In this part, we consider RRE when $p-1<n$ first and then analyze $p>n$ case. In that respect $\hat{B}=I_{p-1}$, which simplifies the upper bound in Theorem 1. We have
@@
for $j=1,\cdots,p$. Note that when $p-1<n$, RRE estimation is equivalent to OLS, and by using (132) of Bunea et al. (2022), $\hat{\alpha}_{RRE,j}=(Y_{-j})^{+}y_{j}=(Y_{-j}^{\prime}Y_{-j})^{-1}Y_{-j}^{\prime}y_{j}.$ We show that after Corollary 2-Remark 4, indeed in an important subcase, OLS will not perform well in finite samples.
Corollary 2.
(RRE: $p-1<n$)
(i). Under Assumptions 1-4, and $Klogn<cn$, with $c>0$ a positive constant, with probability at least $1-1/n^{C_{p}/2}$, and $r_{n}/\delta_{n}\to 1$
$\max_{1\leq j\leq p}\|\hat{\alpha}_{RRE,j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\sigma^{2}\left[\frac{2logn+p}{Cn}\right]+\frac{CK}{\bar{\xi}}\left[1+\frac{1}{\delta_{n}}+\frac{1}{C}\right].$
@@
Remark 4.
  • Note that when $p$ is fixed , $n$ is fixed and $p/n=\gamma\in(0,1)$, where $\gamma$ is a positive fraction between 0 and 1, RRE or equivalently OLS does not perform well in finite samples. This is clear from the first term on the right side in Corollary 2(i). Even when $p/n\to\gamma$, with $p\to\infty,n\to\infty$, there is no consistency.
  • In large samples, a comparison with adaptive PCR shows that since $p>K$ by Assumption 3 we expect adaptive PCR to have a better rate in estimation error. Signal to noise ratio affects both estimators in the same way by reducing the bias term. Note that in the variance term in adaptive PCR, Corollary 1(ii), we may have $\delta_{n}\to\infty$ in the denominator so adaptive PCR has extra advantage of having a better variance than RRE ($p-1<n$). Adaptive PCR will perform better when $p/n=\gamma\in(0,1)$ case unlike RRE here asymptotically. RRE will be consistent when $p/n\to 0$.
  • Note that there is one less bias term in upper bound in RRE compared with adaptive PCR due to $\bar{\Psi}=0$ in RRE. This may be the only advantage in finite sample context.
Now we provide the Corollary for the RRE case when $p-1>n$. First we find it useful to introduce the effective rank as $re(\Sigma_{U,-j}):=\frac{tr(\Sigma_{U,-j})}{\|\Sigma_{U,-j}\|_{2}}.$ Next, we relax Assumption 3(i) that relates $p$ to $n$. We show that in RRE with $p-1>n$ case, we can allow $p$ to grow exponentially in $n$. This implies that we can have much larger, in dimension, precision matrix estimation in RRE compared with adaptive PCR.
Assumption $3^{*}(i)$. We allow $p\leq\frac{exp(C_{p}n/2)}{4}.$ Now we provide our coefficient estimation result that will be helpful for precision matrix estimation based on RRE.
Corollary 3.
(RRE: $p-1>n$)
Under Assumptions 1-2, and Assumptions $3^{*}(i)$-3(ii)-4 and $C>0$, and for all $j=1,\cdots,p$, $re(\Sigma_{U,-j})>Cn,$ for $C$ a positive constant $Klogn<cn$, with $c$ a positive constant, and $\tilde{U}_{-j}$ has independent entries, with $r_{n}/\delta_{n}\to 1$
@@
$\max_{1\leq j\leq p}\|\hat{\alpha}_{RRE,j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\sigma^{2}\left[\frac{2logn+n}{Cn\delta_{n}}\right]+\frac{CK}{\bar{\xi}}\left[1+\frac{1}{\delta_{n}}+\frac{1}{C\delta_{n}}\right].$
@@
  • Note that we may see double descent in RRE for off-diagonal term estimation of the parameter vector $\alpha_{j}^{*}$ in the precision matrix. We fix $n$ and $K$ and analyze the changes in $p$. To see that in the previous Corollary we need $p/n\to 0$ to get estimation errors converging to zero. So very small $p$ in OLS (RRE $p-1<n$) case provides estimation error converging to zero. Then in minimum-norm interpolator case here (RRE with $p-1>n$), with $p/n\to\infty$ with $n$ fixed, and noise $\delta_{n}$ may grow with larger $p$ as in Assumption 2 $\frac{logn+n}{n\delta_{n}}\to 0,$ and $K/\bar{\xi}\to 0$ as discussed above since $r_{n},d_{1n}^{2}$ depend on $p$ and if $d_{1n}^{2}$ dominates $r_{n}$ we may see estimation errors converging to zero, and hence double descent can be seen. To see double descent Assumption $^{*}$(i),(ii) and (8) is helpful.
  • We compare now RRE in this case with adaptive PCR. In RRE there is one less bias term compared with adaptive PCR ($\bar{\Psi}=0$ in RRE) in finite sample upper bound. Since $K<n$, adaptive PCR has a smaller variance term compared with RRE. Asymptotically bias terms have the same rate $K/\bar{\xi}$. So asymptotically adaptive PCR have smaller estimation errors, but in finite samples due to one less bias term in RRE, RRE may have smaller bias, compared to adaptive PCR.
@@ 4.2 Estimation of Diagonal Precision Matrix Entries
We have the following Theorem for the reciprocal of the main $j$-th diagonal term in the precision matrix. The technique in Bing et al. (2021) considers only simple linear regression, while the precision matrix estimation and especially a diagonal term estimation are not handled. The proof of Theorem 2 extends the literature on diagonal term estimation of a large precision matrix. The results and the proof are novel and show the rate of convergence and the non-asymptotic upper bound for estimation errors, as well as the consistency of the estimator. Note that $r_{w1},r_{w2}$ rates are defined just before Assumption 5 in the main text.
Theorem 2.
(i). Under Assumptions 1-4 5(i), with $r_{n}/\delta_{n}\to 1$,
@@
$+$ $O_{p}\left(d_{2n}r_{w2}\right)+O_{p}\left(max(\frac{K^{3/2}}{n^{1/2}},K\frac{\max(\sqrt{logp},\sqrt{logn})}{n^{1/2}})\right).$
(ii). Under Assumptions 1-5, with $r_{n}/\delta_{n}\to 1$,
@@
(iii). Under Assumptions 1-5, with $r_{n}/\delta_{n}\to 1$, where $c>0$ is a positive constant
@@
Remark 6.
  • We show that estimation of the reciprocal of the main diagonal terms in precision matrix estimation depends on number of factors, $K$, sample size-time span $n$, number of assets $p$, and signal-to noise ratio $\bar{\xi}$ as well as three terms $\bar{r},\bar{\Psi},\bar{\eta}$ which represents complexity, prediction errors, and prediction performance respectively. These last three terms are random and will be upper bounded with non-random terms, wpa1, in subcases of RRE and adaptive PCR below in Corollaries A.1-A.3.
@@
  • Number of factors $K$ play a key role, to have Assumption 5(ii), we need $K=o(n^{1/3})$, and there are more restrictions on $K$ in Assumptions 5(i), (iii). Larger number of factors provide a real roadblock in getting consistency.
  • Most important is the fact that estimation of the reciprocal of the main diagonal terms in the precision matrix. This is more difficult compared with estimating each row of the precision matrix (without the main diagonal term). Theorem 1(ii) clearly shows that the rate of estimation error for off-diagonal-row elements is $r_{w2}$, but our Theorem 2 shows for the reciprocal of the main diagonal term, this is more complicated and slower than $r_{w2}$ since via Assumption 5(iii), we need $d_{2n}r_{w2}=o_{p}(1)$ for consistency in addition to other part of Assumption 5(iii) as well as Assumption 5(ii).
  • Note that Theorem 2 excludes the case of RRE with $p-1>n$ case, since formula in (12) is equal to zero due to interpolation. We provide an alternative estimator for RRE with $p-1>n$ case. That case is proved separately in Corollary OA.3. Subcases of adaptive PCR and RRE with $p-1<n$ cases are shown in Corollaries OA.1-OA.2.

### v2

@@ 4 Uniform Consistency of General Estimator
In this section we provide first upper bounds on $l_{2}$ norm errors for each regression coefficient estimate, that is in (10). Next, we consider (11), and obtain upper bounds for the estimation errors.
@@ 4.1 Bounding Estimation Errors: Regression Coefficients
which is the $\hat{r}_{j}$- th largest singular value (squared, scaled by n) of $Y_{-j}P_{\hat{B}}$ ($n\times p-1$ matrix). This last term is obtained when we project $Y_{-j}$ onto range of $\hat{B}$, and will reduce both the bias and variance of estimating the rows of the precision matrix in Theorem 1 below. It can be considered as size of the signal, and its key to consistency of our estimators. Next, we define
@@
Define $\max_{1\leq j\leq p}\hat{r}_{j}:=\bar{r}$, $\min_{1\leq j\leq p}\hat{\eta}_{j}:=\bar{\eta}$, and $\max_{1\leq j\leq p}\hat{\Psi}_{j}:=\bar{\Psi}$. We have the following uniform bounds for estimators for off-diagonal terms in each row of the precision matrix, with $\beta,\Gamma$ as positive large constants, and $\bar{r},\bar{\eta},\bar{\Psi},\bar{\xi}$ are random variables,
Theorem 1
@@
$\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\Gamma\left(\frac{2logn+\bar{r}}{n\bar{\eta}}\right)+\frac{CK}{\bar{\xi}}\left(1+\frac{1}{\beta}+\frac{\bar{\Psi}}{\bar{\eta}\beta}+\frac{\Gamma}{\bar{\eta}\beta}\right).$
(ii) Under Assumptions 1-4,
$\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}^{2}=O_{p}\left(\max(\frac{K}{\bar{\xi}},\frac{K\bar{\Psi}}{\bar{\xi}\bar{\eta}},\frac{K}{\bar{\xi}\bar{\eta}},\frac{logn+\bar{r}}{n\bar{\eta}})\right)=O_{p}(r_{w1}^{2}).$
Remark 1
  • The first term on the right side of Theorem 1(i) is the variance term and the second term is the bias term, and there is no sparsity parameter on the upper bounds and rates.
  • Note that $\bar{\xi}$ plays an essential role in getting consistency. Larger signal-to-noise ratio is needed, especially if the number of factors $K$ grow fast. With using example (8), we see that we cannot have many very weak factors. As an example if the factor loadings are very weak in the following rate $d_{1n}^{2}=O(K)$, there will be no consistency.
  • $\bar{r}$ represents complexity, and affects the $l_{2}$ norm error adversely through the increase of the estimator’s variance. On the other hand, $\bar{\eta}$ is a positive influence, with larger signal improving the performance. It should particularly be noted that this positive influence is reflected through the reduction of both the estimator’s bias and variance. $\bar{\Psi}$ has an adverse effect on $l_{2}$ error. This term increases the bias of the estimator.
@@ 4.1.1 Adaptive PCR
where $(.)^{+}$ is Moore-Penrose inverse of a matrix. Before simplifying the terms in the upper bound in Theorem 1 we define $\hat{\lambda}_{k_{j}}:=\frac{1}{n}\sigma_{k_{j}}^{2}(Y_{-j})$, for example with $k_{j}=1$ we have the largest singular value of $Y_{-j}$, which is $\sigma_{1}^{2}(Y_{-j})/n$ (squared-scaled by $n$). One big issue is the choice of $k_{j}$. To choose $k_{j}$ we adapt the so-called elbow method, (for details, see (19) of Bing et al. (2021))
@@
We show in the proof of Corollary 1 in Appendix
@@
which will simplify the upper bound for $l_{2}$ norm error for our estimators. We prove the following error bounds for adaptive PCR based estimator for the off-diagonal terms in a given row for the precision matrix. Note that three random terms $\bar{r}\leq K$, $\bar{\eta}\geq\beta>0$, and $\bar{\Psi}/\bar{\eta}\leq 1$ with probability approaching one to obtain Corollary 1 from Theorem 1.
Corollary 1
Under Assumptions 1-4 with $0<c^{\prime}<1$ a positive constant, $Klogn\leq cn$, with $c$ a positive constant,
@@
$\max_{1\leq j\leq p}\|\hat{\alpha}_{PCR,j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\Gamma\left[\frac{2logn+K}{n\beta}\right]+\frac{CK}{\bar{\xi}}\left[1+\frac{2}{\beta}+\frac{\Gamma}{\beta^{2}}\right],$
@@
$\max_{1\leq j\leq p}\|\hat{\alpha}_{PCR,j}-\alpha_{j}^{*}\|_{2}^{2}=O_{p}\left(max(\frac{logn+K}{n},\frac{K}{\bar{\xi}})\right).$
Note that noise constant $\Gamma$ play an important role in finite samples by increasing the variance and bias, but $\beta$ has the opposite effect. High dimensional consistency is possible when $p>n$, with $\bar{\xi}$ increasing with $p$ as can be seen in Example in (8). To have a feasible estimation with $K$ estimated, we follow (56) of Bing et al. (2021) or (2.7) of Bing and Wegkamp (2019). Assuming additionally in Corollary 1 that $tr(\Sigma_{U,-j})\geq c^{\prime\prime}\|\Sigma_{U,-j}\|_{2}(n\wedge p)$, with $c^{\prime\prime}>0$ a positive constant, for each $j=1,\cdots,p$, Proposition 8 of Bing et al. (2021) can be extended with a simple union bound and the feasible estimation of the unknown number of factors results do not change Corollary 1.
@@ 4.1.2 Ridgeless Regression Estimation
In this part, we consider RRE when p1<np-1<n first and then analyze $p>n$ case. In that respect $\hat{B}=I_{p-1}$, which simplifies the upper bound in Theorem 1. We have
@@
for $j=1,\cdots,p$. Note that when p1<np-1<n, RRE estimation is equivalent to OLS, and by using (132) of Bunea et al. (2022), $\hat{\alpha}_{RRE,j}=(Y_{-j})^{+}y_{j}=(Y_{-j}^{\prime}Y_{-j})^{-1}Y_{-j}^{\prime}y_{j}.$ Three random terms simplify $\bar{r}<p$, $\bar{\Psi}=0$, and $\bar{\eta}\geq C>0$ with probability approaching one to obtain Corollary 2 from Theorem 1.
Corollary 2
(RRE: p1<np-1<n)
(i). Under Assumptions 1-4, and $Klogn<cn$, with $c>0$ a positive constant, with probability at least $1-1/n^{C_{p}/2}$,
$\max_{1\leq j\leq p}\|\hat{\alpha}_{RRE,j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\Gamma\left[\frac{2logn+p}{Cn}\right]+\frac{CK}{\bar{\xi}}\left[1+\frac{1}{\beta}+\frac{\Gamma}{C\beta}\right].$
@@
Now we provide the Corollary for the RRE case when p1>np-1>n. First we find it useful to introduce the effective rank as $re(\Sigma_{U,-j}):=\frac{tr(\Sigma_{U,-j})}{\|\Sigma_{U,-j}\|_{2}}.$ Next, we relax Assumption 3(i) that relates $p$ to $n$. We show that in RRE with p1>np-1>n case, we can allow $p$ to grow exponentially in $n$. This implies that we can have much larger, in dimension, precision matrix estimation in RRE compared with adaptive PCR.
Assumption $3^{*}(i)$. We allow
$n<p-1\leq\frac{exp(C_{p}n/2)}{4}.$
Now we provide our coefficient estimation result that will be helpful for precision matrix estimation based on RRE. Define $\delta_{n}:=\min_{1\leq j\leq p}tr(\Sigma_{U,-j})/n$. $\delta_{n}$ is nondecreasing in n, since $\Sigma_{U,-j}:p-1\times p-1$ diagonal matrix, with positive elements in the diagonal and $p,n$ grow together under Assumption $3^{*}(i)$ with p1>np-1>n. Note that three random variables in Theorem 1 simplifies in Corollary 3 as $\bar{r}=n$, the bias term is $\bar{\Psi}=0$, and the size of the signal is $\bar{\eta}\geq\delta_{n}$ wpa1.
Corollary 3
(RRE: p1>np-1>n)
Under Assumptions 1-2, and Assumptions $3^{*}(i)$-3(ii)-4 and $C>0$, and for all $j=1,\cdots,p$, $re(\Sigma_{U,-j})>Cn,$ for $C$ a positive constant $Klogn<cn$, with $c$ a positive constant, and $\tilde{U}_{-j}$ has independent entries,
@@
$\max_{1\leq j\leq p}\|\hat{\alpha}_{RRE,j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\Gamma\left[\frac{2logn+n}{Cn\delta_{n}}\right]+\frac{CK}{\bar{\xi}}\left[1+\frac{1}{\beta}+\frac{\Gamma}{\beta\delta_{n}}\right].$
@@
We compare now RRE in this case with adaptive PCR. In RRE there is one less bias term compared with adaptive PCR ($\bar{\Psi}=0$ in RRE) in finite sample upper bound. Since $K<n$, adaptive PCR has a smaller variance term compared with RRE. Asymptotically bias terms have the same rate $K/\bar{\xi}$.
@@ 4.2 Estimation of Diagonal Precision Matrix Entries
We have the following Theorem for the reciprocal of the main $j$-th diagonal term in the precision matrix. The technique in Bing et al. (2021) considers only simple linear regression, while the precision matrix estimation and especially a diagonal term estimation are not handled. The proof of Theorem 2 extends the literature on diagonal term estimation of a large precision matrix. The results and the proof are novel and show the rate of convergence and the non-asymptotic upper bound for estimation errors, as well as the consistency of the estimator. Note that $r_{w1}$ rate is defined just before Assumption 5 in the main text.
Theorem 2
(i). Under Assumptions 1-4 5(i),
@@
$+$ $O_{p}\left(d_{2n}r_{w1}\right)+O_{p}\left(max(\frac{K^{3/2}}{n^{1/2}},K\frac{\max(\sqrt{logp},\sqrt{logn})}{n^{1/2}})\right).$
(ii). Under Assumptions 1-5,
@@
(iii). Under Assumptions 1-5, where $c>0$ is a positive constant
@@
Remark 2
@@
  • Theorem 1(ii) clearly shows that the rate of estimation error for off-diagonal-row elements is $r_{w1}$, but our Theorem 2 shows for the reciprocal of the main diagonal term, this is slower than $r_{w1}$ since via Assumption 5(iii), we need $d_{2n}r_{w1}=o_{p}(1)$ for consistency in addition to other part of Assumption 5(iii) as well as Assumption 5(ii).
  • Note that Theorem 2 excludes the case of RRE with p1>np-1>n case, since formula in (11) is equal to zero due to interpolation. We provide an alternative estimator for RRE with p1>np-1>n case in Corollary A.3 in Online Appendix.